# Compressed and Smooth Latent Space for Text Diffusion Modeling

### v2

@@ 1 Introduction
Autoregressive (AR) language models are a de facto gold standard for text generation (radford2018improving, ; radford2019language, ; brown2020gpt3, ). By factorizing the probability of a sequence into a product of conditional token probabilities, they transform the global generation task into a series of local next-token prediction tasks that can be optimized efficiently via teacher forcing (williams1989learning, ).
The same sequential factorization, however, creates several structural bottlenecks. First, decoding is inherently sequential: the time to produce a sequence grows linearly with its length because each token must await the completion of its predecessor. Second, the process suffers from exposure bias: a single early mistake contaminates the context for all subsequent predictions and cannot be corrected without restarting generation (ranzato2015sequence, ; bengio2015scheduled, ; he2019exposure, ). Third, maximizing local log-likelihood encourages the model to privilege fluency over factual accuracy, resulting in the well-documented phenomenon of hallucination (maynez2020faithfulness, ; ji2023survey, ). Finally, because decisions are made token-by-token, the model lacks an explicit global plan and therefore struggles to maintain long-range logical or narrative coherence (ye2024beyond, ).
In stark contrast, computer vision has been reshaped by diffusion models (ho2020ddpm, ; nichol2021improved, ), particularly by latent diffusion (rombach2022ldm, ; blattmann2023stable, ; andreas2023, ), which first compresses a high-resolution image or video into a compact latent representation and then performs the diffusion process in that space. Operating in a low-dimensional latent manifold reduces computational cost by orders of magnitude and enables breathtaking advances in image and video synthesis. Yet recent works (kouzelis2025eq, ; skorokhodov2025improving, ) also show that diffusion is highly sensitive to the geometry of the latent space: poorly designed latent representations can destabilize training and degrade sample quality, underscoring the need for principled mechanisms to construct diffusable [^1] Diffusable representations refer to latent spaces that permit effective modeling by a diffusion process. representations.
@@ 2 Related work
Early efforts to bring latent diffusion models into natural language generation treat the diffusion module mainly as a pre-processor that supplies conditioning vectors for an autoregressive decoder. LD4LG (lovelace2023latent, ) trains a diffusion model on compressed hidden states from BART, then feeds the generated latent vectors into the BART decoder to produce text. PLANNER (zhang2023planner, ) adopts a similar two-stage recipe: a fine-tuned BERT encoder produces a 16-token variational latent code, a diffusion model refines that code, and a GPT-2 decoder generates text from this code. While both systems improve controllability, their heavy reliance on powerful autoregressive decoders makes it difficult to isolate and evaluate the intrinsic generative capacity of the latent-space diffusion itself.
A complementary line of work applies diffusion directly to the continuous embeddings of pretrained encoders. TEncDM (shabalin2025tencdm, ) demonstrates that full-length BERT representations can be modeled with Gaussian diffusion and decoded into coherent text without an intermediate autoregressive step. Crucially, however, TEncDM retains the original sequence dimensionality, leaving open a question: How aggressively can such representations be compressed before generation quality collapses?
@@ 3 Preliminary
Diffusion models (ho2020ddpm, ; nichol2021improved, ) learn a data distribution by reversing a progressive noising process. Given a clean sample $z_{0}\sim p_{\text{data}}$, the forward dynamic corrupts the input with Gaussian noise whose magnitude is controlled by a continuous time index $t\in[0,1]$:
@@ 3.1 Latent diffusion
Latent Diffusion Models (LDMs) (rombach2022ldm, ; andreas2023, ) improve both compute efficiency and quality by learning the diffusion process in a compact latent space rather than in pixel or token space. An autoencoder with encoder $E$ and decoder $D$ is first trained to achieve high-fidelity reconstruction, $\mathbf{\hat{w}}~=~D\!\bigl(E(\mathbf{w})\bigr)~\approx~\mathbf{w},$ where $\mathbf{z}=E(\mathbf{w})$ denotes the low-dimensional latent. Eqs. (1)–(2) yield a diffusion model trained in latent space whose network can be shallower, and faster than a counterpart operating in the original space, yet still achieves comparable perceptual quality after decoding $D(\mathbf{z})$.
@@ 4.1 Overview
We initialise the pipeline with a pretrained contextual encoder $E_{\text{text}}$ (BERT-base (devlin2019bert, ) by default) that remains frozen during all subsequent training stages. For an input sequence of $L$ tokens $\mathbf{w}=(w_{1},\dots,w_{L})$, the encoder produces a matrix of hidden states $\mathbf{h}=E_{\text{text}}(\mathbf{w})\in~\mathbb{R}^{L\times d}$, where $d=768$. Each row supplies a semantically rich representation of its corresponding token tenney-etal-2019-bert . These representations provide a high-fidelity starting point for the compression stage.
@@ 5 Latent‑space properties that facilitate diffusion training
This section embarks on an exploration of the text‑latent manifold, asking which of its intrinsic features govern the performance of a diffusion model. All experiments reported here employ autoencoders that compress 128‑token texts into 16 latent vectors. The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset. Across all experiments, we generate 1 000 samples and report each metric as the average over these samples. Our experiments spotlight two indispensable qualities — manifold smoothness and resilience to perturbations. The remainder of the section introduces an analysis that quantifies both attributes.
@@ 5.1 Smoothness of the latent manifold
Once the autoencoder has been trained, each text $\mathbf{w}$ corresponds to a set of latent vectors $\mathbf{z}$ that can be decoded back into the original text, $\mathbf{w}=D(\mathbf{z})$. Nonetheless, the latent space is far from fully charted: broad regions contain vectors to which no text has ever been assigned, leaving the behaviour of the data density there unknown. Because our diffusion model captures the distribution of texts via the distribution of these latents, yet observes only a finite subset during training, its ability to generalize depends critically on how smoothly that density varies across the manifold. Put differently, the degree to which a locally estimated score function extends to unexplored territory is dictated by manifold smoothness. Inspired by PLANNER (zhang2023planner, ), we investigate this property by conducting the following experiment, which simultaneously evaluates performance at points the model does not encounter during training.
@@ 5.2 Reducing the train–inference mismatch
To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$. Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout. Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@ 6.1 Empirical analysis of autoencoder training regimes
In this part of our study we assess the contribution of each proposed latent space augmentation to the final quality of the diffusion model. For the ablation study, we compress text from 128 BERT representations down to 16 latent vectors. All autoencoders are trained on the Wikipedia dataset, while the diffusion models are trained on the ROCStories (mostafazadeh2016corpus, ) dataset.
@@ Results.
Table˜2 demonstrates that each proposed feature contributes significantly to the model performance. Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%. Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations. This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full‑length, uncompressed BERT representations (0.767 vs. 0.762 for TEncDM (shabalin2025tencdm, )). Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
@@
Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset. Detailed results of this evaluation are presented in LABEL:app:ablation_wikipedia. Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section˜D.1.
@@ 7 Comparison across generative paradigms
We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )). For unconditional story generation we employ metrics discussed in Section˜6.1. For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ). The detoxification task is evaluated using BLEU-4 (BLEU). We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens. Additional metrics and implementation details are provided in the Section˜D.2.
@@
We benchmark our proposed method against a diverse set of generative paradigms. These include Gaussian diffusion on embeddings (DiffuSeq (gong2022diffuseq, ), SeqDiffuSeq (yuan2022seqdiffuseq, ), AR-Diffusion (wu2023ar, )), simplex-based diffusion (TESS (mahabadi2023tess, )), masked diffusion (SEDD (lou2023discrete, )), alternative latent diffusion baselines (LD4LG (lovelace2023latent, ), TEncDM (shabalin2025tencdm, )), and autoregressive models (GPT-2 (radford2019language, ) and GPT-Neo (gpt-neo, )). All models operate within a comparable parameter scale, ranging from approximately 100M to 200M. To ensure fairness, all baselines were faithfully reimplemented and trained on generative tasks using training protocols closely aligned with those described by their original authors.
@@ 8 Experiments on OpenWebText
In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb, ) dataset.
@@ 8.2 Comparison on Long-Sequence Generation
We further assess our model’s capabilities on the challenging task of generating 512-token sequences. We compare our model, Cosmos, in two configurations: without compression ($N=512$ latent vectors) and with 2x compression ($N=256$). These are benchmarked against TEncDM (shabalin2025tencdm, ), a latent diffusion baseline, and GIDD (von2025generalized, ), a strong discrete diffusion model. The results, presented in Table 6, offer several key insights into the advantages of our approach.
@@ Appendix A Dataset descriptions
This dataset [24] consists of five-sentence stories and serves as a well-established, small-scale, unconditional benchmark in language diffusion research. It contains a total of 98,161 instances, of which 88,161 are used for training, 10,000 for validation.
@@
For large-scale experiments, we use the English Wikipedia subset from the ROOTS corpus [15]. The resulting dataset comprises 8.6 million sequences, with 50,000 held out for validation.
@@
This dataset [26] is used for abstractive summarization and comprises 204,000 BBC news articles paired with summaries covering diverse topics (e.g., sports, politics). It includes 204,045 training instances, 11,332 validation instances, and 11,334 test instances. We truncate input articles to 512 tokens and limit reference summaries to 50 tokens.
@@
This question-answering dataset [31] requires identifying the text span that answers a given question or indicating that no answer is possible from the provided context. For generative modeling, we reverse the task: given a context and an answer, the model generates the corresponding question. The dataset contains 130,319 training and 11,873 test instances. Contexts are truncated to 512 tokens.
@@
For small-scale conditional generation experiments, we use the ParaDetox dataset [17], which comprises 19,766 pairs of toxic and neutral comments. Both context and target sequences are truncated to 40 tokens.
@@ B.1 Gaussian Diffusion on Embeddings
One line of work focuses on applying Gaussian diffusion directly to word embeddings. Models like DiffuSeq [10] and SeqDiffuSeq [45] are designed for sequence-to-sequence text generation tasks. DiffuSeq employs a partial noising strategy and is trained end-to-end in a classifier-free manner. SeqDiffuSeq utilizes an encoder-decoder Transformer architecture and incorporates self-conditioning and an adaptive noise schedule to enhance performance. AR-Diffusion [43] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left. This is achieved by using a dynamic number of denoising steps that varies with the token’s position.
@@ B.2 Simplex-Based Diffusion
Another direction explores alternative spaces for the diffusion process. TESS [20], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space. This approach, combined with a novel self-conditioning technique, demonstrates strong performance on various natural language generation tasks while requiring fewer diffusion steps.

### v3

# Cosmos: Compressed and Smooth Latent Space for Text Diffusion Modeling

@@ 1 Introduction
Autoregressive (AR) language models are a de facto gold standard for text generation (radford2018improving; radford2019language; brown2020gpt3). By factorizing the probability of a sequence into a product of conditional token probabilities, they transform the global generation task into a series of local next-token prediction tasks that can be optimized efficiently via teacher forcing (williams1989learning).
The same sequential factorization, however, creates several structural bottlenecks. First, decoding is inherently sequential: the time to produce a sequence grows linearly with its length because each token must await the completion of its predecessor. Second, the process suffers from exposure bias: a single early mistake contaminates the context for all subsequent predictions and cannot be corrected without restarting generation (ranzato2015sequence; bengio2015scheduled; he2019exposure). Third, maximizing local log-likelihood encourages the model to privilege fluency over factual accuracy, resulting in the well-documented phenomenon of hallucination (maynez2020faithfulness; ji2023survey). Finally, because decisions are made token-by-token, the model lacks an explicit global plan and therefore struggles to maintain long-range logical or narrative coherence (ye2024beyond).
In stark contrast, computer vision has been reshaped by diffusion models (ho2020ddpm; nichol2021improved), particularly by latent diffusion (rombach2022ldm; blattmann2023stable; andreas2023), which first compresses a high-resolution image or video into a compact latent representation and then performs the diffusion process in that space. Operating in a low-dimensional latent manifold reduces computational cost by orders of magnitude and enables breathtaking advances in image and video synthesis. Yet recent works (kouzelis2025eq; skorokhodov2025improving) also show that diffusion is highly sensitive to the geometry of the latent space: poorly designed latent representations can destabilize training and degrade sample quality, underscoring the need for principled mechanisms to construct diffusable [^1] Diffusable representations refer to latent spaces that permit effective modeling by a diffusion process. representations.
@@ 2 Related work
Early efforts to bring latent diffusion models into natural language generation treat the diffusion module mainly as a pre-processor that supplies conditioning vectors for an autoregressive decoder. LD4LG (lovelace2023latent) trains a diffusion model on compressed hidden states from BART, then feeds the generated latent vectors into the BART decoder to produce text. PLANNER (zhang2023planner) adopts a similar two-stage recipe: a fine-tuned BERT encoder produces a 16-token variational latent code, a diffusion model refines that code, and a GPT-2 decoder generates text from this code. While both systems improve controllability, their heavy reliance on powerful autoregressive decoders makes it difficult to isolate and evaluate the intrinsic generative capacity of the latent-space diffusion itself.
A complementary line of work applies diffusion directly to the continuous embeddings of pretrained encoders. TEncDM (shabalin2025tencdm) demonstrates that full-length BERT representations can be modeled with Gaussian diffusion and decoded into coherent text without an intermediate autoregressive step. Crucially, however, TEncDM retains the original sequence dimensionality, leaving open a question: How aggressively can such representations be compressed before generation quality collapses?
@@ 3 Preliminary
Diffusion models (ho2020ddpm; nichol2021improved) learn a data distribution by reversing a progressive noising process. Given a clean sample $z_{0}\sim p_{\text{data}}$, the forward dynamic corrupts the input with Gaussian noise whose magnitude is controlled by a continuous time index $t\in[0,1]$:
@@ 3.1 Latent diffusion
Latent Diffusion Models (LDMs) (rombach2022ldm; andreas2023) improve both compute efficiency and quality by learning the diffusion process in a compact latent space rather than in pixel or token space. An autoencoder with encoder $E$ and decoder $D$ is first trained to achieve high-fidelity reconstruction, $\mathbf{\hat{w}}~=~D\!\bigl(E(\mathbf{w})\bigr)~\approx~\mathbf{w},$ where $\mathbf{z}=E(\mathbf{w})$ denotes the low-dimensional latent. Eqs. (1)–(2) yield a diffusion model trained in latent space whose network can be shallower, and faster than a counterpart operating in the original space, yet still achieves comparable perceptual quality after decoding $D(\mathbf{z})$.
@@ 4.1 Overview
We initialise the pipeline with a pretrained contextual encoder $E_{\text{text}}$ (BERT-base (devlin2019bert) by default) that remains frozen during all subsequent training stages. For an input sequence of $L$ tokens $\mathbf{w}=(w_{1},\dots,w_{L})$, the encoder produces a matrix of hidden states $\mathbf{h}=E_{\text{text}}(\mathbf{w})\in~\mathbb{R}^{L\times d}$, where $d=768$. Each row supplies a semantically rich representation of its corresponding token tenney-etal-2019-bert. These representations provide a high-fidelity starting point for the compression stage.
@@ 5 Latent‑space properties that facilitate diffusion training
This section embarks on an exploration of the text‑latent manifold, asking which of its intrinsic features govern the performance of a diffusion model. All experiments reported here employ autoencoders that compress 128‑token texts into 16 latent vectors. The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus) dataset. Across all experiments, we generate 1 000 samples and report each metric as the average over these samples. Our experiments spotlight two indispensable qualities — manifold smoothness and resilience to perturbations. The remainder of the section introduces an analysis that quantifies both attributes.
@@ 5.1 Smoothness of the latent manifold
Once the autoencoder has been trained, each text $\mathbf{w}$ corresponds to a set of latent vectors $\mathbf{z}$ that can be decoded back into the original text, $\mathbf{w}=D(\mathbf{z})$. Nonetheless, the latent space is far from fully charted: broad regions contain vectors to which no text has ever been assigned, leaving the behaviour of the data density there unknown. Because our diffusion model captures the distribution of texts via the distribution of these latents, yet observes only a finite subset during training, its ability to generalize depends critically on how smoothly that density varies across the manifold. Put differently, the degree to which a locally estimated score function extends to unexplored territory is dictated by manifold smoothness. Inspired by PLANNER (zhang2023planner), we investigate this property by conducting the following experiment, which simultaneously evaluates performance at points the model does not encounter during training.
@@ 5.2 Reducing the train–inference mismatch
To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$. Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout. Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@ 6.1 Empirical analysis of autoencoder training regimes
In this part of our study we assess the contribution of each proposed latent space augmentation to the final quality of the diffusion model. For the ablation study, we compress text from 128 BERT representations down to 16 latent vectors. All autoencoders are trained on the Wikipedia dataset, while the diffusion models are trained on the ROCStories (mostafazadeh2016corpus) dataset.
@@ Results.
Table˜2 demonstrates that each proposed feature contributes significantly to the model performance. Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%. Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations. This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full‑length, uncompressed BERT representations (0.767 vs. 0.762 for TEncDM (shabalin2025tencdm)). Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
@@
Additionally, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section˜D.1.
@@ 7 Comparison across generative paradigms
We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus)), text summarization (XSum (narayan2018don)), detoxification (ParaDetox (logacheva-etal-2022-paradetox)), and question generation (SQuAD2.0 (minaee2021deep)). For unconditional story generation we employ metrics discussed in Section˜6.1. For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore). The detoxification task is evaluated using BLEU-4 (BLEU). We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens. Additional metrics and implementation details are provided in the Section˜D.2.
@@
We benchmark our proposed method against a diverse set of generative paradigms. These include Gaussian diffusion on embeddings (DiffuSeq (gong2022diffuseq), SeqDiffuSeq (yuan2022seqdiffuseq), AR-Diffusion (wu2023ar)), simplex-based diffusion (TESS (mahabadi2023tess)), masked diffusion (SEDD (lou2023discrete)), alternative latent diffusion baselines (LD4LG (lovelace2023latent), TEncDM (shabalin2025tencdm)), and autoregressive models (GPT-2 (radford2019language) and GPT-Neo (gpt-neo)). All models operate within a comparable parameter scale, ranging from approximately 100M to 200M. To ensure fairness, all baselines were faithfully reimplemented and trained on generative tasks using training protocols closely aligned with those described by their original authors.
@@ 8 Experiments on OpenWebText
In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb) dataset.
@@ 8.2 Comparison on Long-Sequence Generation
We further assess our model’s capabilities on the challenging task of generating 512-token sequences. We compare our model, Cosmos, in two configurations: without compression ($N=512$ latent vectors) and with 2x compression ($N=256$). These are benchmarked against TEncDM (shabalin2025tencdm), a latent diffusion baseline, and GIDD (von2025generalized), a strong discrete diffusion model. The results, presented in Table 6, offer several key insights into the advantages of our approach.
@@ Appendix A Dataset descriptions
This dataset [rocstories] consists of five-sentence stories and serves as a well-established, small-scale, unconditional benchmark in language diffusion research. It contains a total of 98,161 instances, of which 88,161 are used for training, 10,000 for validation.
@@
For large-scale experiments, we use the English Wikipedia subset from the ROOTS corpus [laurenccon2022bigscience]. The resulting dataset comprises 8.6 million sequences, with 50,000 held out for validation.
@@
This dataset [xsum-emnlp] is used for abstractive summarization and comprises 204,000 BBC news articles paired with summaries covering diverse topics (e.g., sports, politics). It includes 204,045 training instances, 11,332 validation instances, and 11,334 test instances. We truncate input articles to 512 tokens and limit reference summaries to 50 tokens.
@@
This question-answering dataset [rajpurkar-etal-2018-know] requires identifying the text span that answers a given question or indicating that no answer is possible from the provided context. For generative modeling, we reverse the task: given a context and an answer, the model generates the corresponding question. The dataset contains 130,319 training and 11,873 test instances. Contexts are truncated to 512 tokens.
@@
For small-scale conditional generation experiments, we use the ParaDetox dataset [logacheva-etal-2022-paradetox], which comprises 19,766 pairs of toxic and neutral comments. Both context and target sequences are truncated to 40 tokens.
@@ B.1 Gaussian Diffusion on Embeddings
One line of work focuses on applying Gaussian diffusion directly to word embeddings. Models like DiffuSeq [gong2022diffuseq] and SeqDiffuSeq [yuan2022seqdiffuseq] are designed for sequence-to-sequence text generation tasks. DiffuSeq employs a partial noising strategy and is trained end-to-end in a classifier-free manner. SeqDiffuSeq utilizes an encoder-decoder Transformer architecture and incorporates self-conditioning and an adaptive noise schedule to enhance performance. AR-Diffusion [wu2023ar] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left. This is achieved by using a dynamic number of denoising steps that varies with the token’s position.
@@ B.2 Simplex-Based Diffusion
Another direction explores alternative spaces for the diffusion process. TESS [mahabadi2023tess], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space. This approach, combined with a novel self-conditioning technique, demonstrates strong performance on various natural language generation tasks while requiring fewer diffusion steps.